# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ B.1 Crucial tools
Let $Q=\{Q_{i}\}_{i=1}^{m}$ be i.i.d. $d$-dimensional random vectors from a distribution $F$ on ${\bf R}^{d}$, where $Q_{i}=(Q_{i,1},\dots,Q_{i,d})$. Let $\hat{Q}=\{\hat{Q}_{i}\}_{i=1}^{m}$ be a perturbed version of $Q$, which can be potentially dependent. Let $K:{\bf R}^{d-1}\to[0,\infty)$ be a bounded (by a constant $K_{\max}$) and Lipschitz continuous kernel with $\int_{{\bf R}^{d-1}}K(u)\,du=1$. Let $h>0$ be a bandwidth parameter. Define the kernel-based conditional CDF estimator:
$\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\,\big{|}\,u)\;:=\;\frac{\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}\hat{Q}_{i,d}\leq x\bigr{\}}\;K\!\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)}}{\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)}},$
@@
  • The marginal density of the conditioning variables, $f_{1:(d-1)}(u)$, is continuous and bounded. Furthermore, there exists a compact set $\mathcal{U}\subset{\bf R}^{d-1}$ such that $\inf_{u\in\mathcal{U}}f_{1:(d-1)}(u)>0$.
  • The conditional CDF $F^{\,d\mid 1:(d-1)}(x\mid u)$ is Lipschitz continuous with respect to $u=(x_{1},\dots,x_{d-1})$.
@@
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\Bigl{|}\,\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\;-\;F^{\,d\mid 1:(d-1)}(x\mid u)\Bigr{|}\;=\;o_{p}(1).$
Hence $\widehat{F}_{m}^{\,d\mid 1:(d-1)}$ converges uniformly to $F^{\,d\mid 1:(d-1)}$ in probability over $\mathcal{U}\times{\bf R}$.
@@
The perturbation control strategy we use in this proof is very similar to the control of indicator functions in the proof of Lemma B.1. The main difference is that we have additional smoother kernels (which are indeed easier to control).
Define
$\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid x_{1},\dots,x_{d-1})$
to be the same kernel estimator but using $\{Q_{i}\}$, i.e.,
$\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid x_{1},\dots,x_{d-1})\;=\;\frac{\sum_{i=1}^{m}\mathbf{1}\{Q_{i,d}\leq x\}\;K\Bigl{(}\tfrac{x_{1}-Q_{i,1}}{h},\dots,\tfrac{x_{d-1}-Q_{i,d-1}}{h}\Bigr{)}}{\sum_{i=1}^{m}K\Bigl{(}\tfrac{x_{1}-Q_{i,1}}{h},\dots,\tfrac{x_{d-1}-Q_{i,d-1}}{h}\Bigr{)}}.$
Then for any $(x_{1},\dots,x_{d-1},x)$,
$\bigl{|}\widehat{F}_{m}^{\,d\mid 1:(d-1)}-F^{\,d\mid 1:(d-1)}\bigr{|}\;\leq\;\underbrace{\bigl{|}\widehat{F}_{m}^{\,d\mid 1:(d-1)}-\widetilde{F}_{m}^{\,d\mid 1:(d-1)}\bigr{|}}_{(\text{I})}\;+\;\underbrace{\bigl{|}\widetilde{F}_{m}^{\,d\mid 1:(d-1)}-F^{\,d\mid 1:(d-1)}\bigr{|}}_{(\text{II})}.$
Next, we discuss the convergence bound for the two terms separately.
Term (II):
Term $(\text{II})$ represents the uniform error of a standard kernel-based conditional CDF estimator. Under the assumptions that $f_{1:(d-1)}(u)$ is continuous and bounded below by a positive constant on $\mathcal{U}$, and $F^{\,d\mid 1:(d-1)}(x\mid u)$ is Lipschitz continuous with respect to $u$, along with the chosen bandwidth $h$ such that $h\to 0$ and $mh^{d-1}\to\infty$, standard nonparametric estimation results (e.g., from Wasserman [2006] for rates of kernel regression) give the following uniform convergence bound:
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\bigl{|}\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)-F^{\,d\mid 1:(d-1)}(x\mid u)\bigr{|}\;=\;O_{P}\!\Bigl{(}(m\,h^{\,d-1})^{-\tfrac{1}{2}}+h\Bigr{)}.$
Term (I):
We now compare the estimator with perturbed data $\{\hat{Q}_{i}\}$ to the same estimator with unperturbed data $\{Q_{i}\}$. For any fixed $(x,u)$, write
$\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\;=\;\frac{A(x,u)}{B(u)},\quad\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\;=\;\frac{C(x,u)}{D(u)},$
@@
$A(x,u)\;=\;\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}\hat{Q}_{i,d}\leq x\bigr{\}}\,K\!\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)},\quad B(u)\;=\;\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)},$
$C(x,u)\;=\;\sum_{i=1}^{m}\mathbf{1}\!\bigl{\{}Q_{i,d}\leq x\bigr{\}}\,K\!\Bigl{(}\tfrac{u-Q_{i,1:(d-1)}}{h}\Bigr{)},\quad D(u)\;=\;\sum_{i=1}^{m}K\!\Bigl{(}\tfrac{u-Q_{i,1:(d-1)}}{h}\Bigr{)}.$
Their difference is
$\frac{A(x,u)}{B(u)}-\frac{C(x,u)}{D(u)}\;=\;\frac{\,A(x,u)D(u)-C(x,u)B(u)\,}{\,B(u)D(u)\,}.$
We will show that the numerator |A(x,u)D(u)C(x,u)B(u)||A(x,u)D(u)-C(x,u)B(u)| is of order $O_{P}(m(\delta_{m}+\delta_{m}/h)+\sqrt{m})$ and $B(u),\,D(u)=\Omega_{P}(m)$, giving
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\Bigl{|}\frac{A(x,u)}{B(u)}-\frac{C(x,u)}{D(u)}\Bigr{|}\;=\;O_{P}\!\Bigl{(}\delta_{m}+\frac{1}{\sqrt{m}}+\frac{\delta_{m}}{h}\Bigr{)}.$
We analyze the terms:
  • Denominator Control: For $u\in\mathcal{U}$, we are given that $\inf_{u\in\mathcal{U}}f_{1:(d-1)}(u)>0$. By the uniform consistency of kernel density estimators (under the continuity of $f_{1:(d-1)}$, boundedness of $K$, and $h\to 0$, $mh^{d-1}\to\infty$), we have that $D(u)/m\to f_{1:(d-1)}(u)$ uniformly in probability over $\mathcal{U}$. Thus, $D(u)=\Omega_{P}(m)$ uniformly over $\mathcal{U}$. Similarly, $B(u)=\Omega_{P}(m)$ uniformly over $\mathcal{U}$ (as the perturbation $\delta_{m}$ is small, it does not change the order of the sum). Therefore, $B(u)D(u)=\Omega_{P}(m^{2})$.
  • Numerator Control: We write the numerator as:
    A(x,u)D(u)C(x,u)B(u)=(A(x,u)C(x,u))D(u)C(x,u)(B(u)D(u)).A(x,u)D(u)-C(x,u)B(u)=(A(x,u)-C(x,u))D(u)-C(x,u)(B(u)-D(u)).
    We need to bound each part of this expression uniformly.
    $|B(u)-D(u)|=\left|\sum_{i=1}^{m}\left[K\Bigl{(}\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\Bigr{)}-K\Bigl{(}\tfrac{u-Q_{i,1:(d-1)}}{h}\Bigr{)}\right]\right|.$
    Since $K$ is Lipschitz continuous with constant $L_{K}$, and $\|\hat{Q}_{i}-Q_{i}\|\leq\delta_{m}$: Each summand is bounded by $L_{K}\frac{\|\hat{Q}_{i,1:(d-1)}-Q_{i,1:(d-1)}\|}{h}\leq L_{K}\frac{\|\hat{Q}_{i}-Q_{i}\|}{h}\leq L_{K}\frac{\delta_{m}}{h}$. Summing over $m$ terms, we get:
    $\sup_{u\in\mathcal{U}}|B(u)-D(u)|\leq mL_{K}\frac{\delta_{m}}{h}=O(m\delta_{m}/h).$
    This is a deterministic uniform bound.
    $A(x,u)-C(x,u)=\sum_{i=1}^{m}\left[\mathbf{1}\{\hat{Q}_{i,d}\leq x\}K(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h})-\mathbf{1}\{Q_{i,d}\leq x\}K(\tfrac{u-Q_{i,1:(d-1)}}{h})\right].$
    Let $\Delta_{i}(x,u)$ be the $i$-th summand. We use the triangle inequality:
    $|\Delta_{i}(x,u)|$ $\leq$ $\underbrace{\bigl{|}\mathbf{1}\{\hat{Q}_{i,d}\leq x\}-\mathbf{1}\{Q_{i,d}\leq x\}\bigr{|}\,K(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h})}_{\text{indicator contribution}}$
    $\;+\;\underbrace{\mathbf{1}\{Q_{i,d}\leq x\}\,\bigl{|}K(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h})-K(\tfrac{u-Q_{i,1:(d-1)}}{h})\bigr{|}}_{\text{kernel contribution}}.$
    For the kernel part: Each term is bounded by $1\cdot L_{K}\frac{\delta_{m}}{h}$. Summing over $m$ terms, this part contributes $O(m\delta_{m}/h)$ to $\sup_{u,x}|A(x,u)-C(x,u)|$.
    For the indicator part: Since $K(\cdot)\leq K_{\max}$, we have
    $\sum_{i=1}^{m}\bigl{|}\mathbf{1}\{\hat{Q}_{i,d}\leq x\}-\mathbf{1}\{Q_{i,d}\leq x\}\bigr{|}\,K(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h})\leq K_{\max}\sum_{i=1}^{m}\bigl{|}\mathbf{1}\{\hat{Q}_{i,d}\leq x\}-\mathbf{1}\{Q_{i,d}\leq x\}\bigr{|}.$
    Let $D_{i}(x)=\bigl{|}\mathbf{1}\{\hat{Q}_{i,d}\leq x\}-\mathbf{1}\{Q_{i,d}\leq x\}\bigr{|}$. The sum $\sum_{i=1}^{m}D_{i}(x)$ is a random variable that depends only on $x$ and the $d$-th components of the data. From Lemma B.1 (univariate case, taking $d=1$ and $p=1$), we know that and since $\sup_{x\in{\bf R}}E[D_{i}(x)]\leq C_{\mathrm{up}}\cdot(2\delta_{m})=O(\delta_{m})$, we have
    $\sup_{x\in{\bf R}}\left|\sum_{i=1}^{m}D_{i}(x)\right|=O(m\delta_{m})+O_{P}(\sqrt{m}).$
    So the indicator part is bounded by $O_{P}(m\delta_{m}+\sqrt{m})$.
    Combining the two parts, we have
    $\sup_{u\in\mathcal{U},\,x\in{\bf R}}|A(x,u)-C(x,u)|=O_{P}(m\delta_{m}+\sqrt{m}+m\delta_{m}/h).$
    Note that we can similalry see $\sup_{u,x}|C(x,u)|=O_{P}(m)$. Therefore, we get
    $\sup_{u\in\mathcal{U},\,x\in{\bf R}}|A(x,u)D(u)-C(x,u)B(u)|$
    $\leq$ $\sup_{u,x}|A(x,u)-C(x,u)|\sup_{u}|D(u)|+\sup_{u,x}|C(x,u)|\sup_{u}|B(u)-D(u)|$
    $=$ $O_{P}(m^{2}\delta_{m}+m\sqrt{m}+m^{2}\delta_{m}/h).$
Combining the bounds for the numerator and denominator:
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\Bigl{|}\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)-\widetilde{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\Bigr{|}=O_{P}\left(\frac{m^{2}\delta_{m}+m\sqrt{m}+m^{2}\delta_{m}/h}{m^{2}}\right)$
@@
$O_{P}\left(\delta_{m}+\frac{\sqrt{m}}{m}+\frac{\delta_{m}}{h}\right)=O_{P}\left(\delta_{m}+\frac{1}{\sqrt{m}}+\frac{\delta_{m}}{h}\right).$
Given the conditions $\delta_{m}=o(h)$ and $\delta_{m}\to 0$, and $1/\sqrt{m}\to 0$, this term vanishes in probability, i.e., $o_{P}(1)$.
Combining the vanishing rates of Term (I) and Term (II):
@@
$\sup_{u\in\mathcal{U},\,x\in{\bf R}}\Bigl{|}\,\widehat{F}_{m}^{\,d\mid 1:(d-1)}(x\mid u)\;-\;F^{\,d\mid 1:(d-1)}(x\mid u)\Bigr{|}\;=\;O_{P}\!\Bigl{(}(m\,h^{\,d-1})^{-\tfrac{1}{2}}+h\Bigr{)}+O_{P}\left(\delta_{m}+\frac{1}{\sqrt{m}}+\frac{\delta_{m}}{h}\right).$
@@
Let $F$ be a continuous and strictly increasing cumulative distribution function (CDF) (or conditional CDF) on $\bf R$}, and let $F^{-1}$ denote its inverse function. Assume the conditions of Corollary B.1 hold. Let $\hat{F}_{m}$ be the smoothed CDF estimator of Corollary B.1. Then, for any closed interval $a,b]\subset(0,1)$, the inverse functions $\hat{F}_{m}^{-1}$ converge uniformly in probability to $F^{-1}$ on $[a,b]$; that is,
$\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
@@
$P\left(\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|\geq\varepsilon\right)<\eta.$
@@
$P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|>\varepsilon\right)\xrightarrow{m\to\infty}0.$
@@
$|u-v|<\delta\implies|F^{-1}(u)-F^{-1}(v)|<\frac{\varepsilon}{2}.$
@@
$\mathscr{A}_{m}=\left\{\sup_{x\in{\bf R}}|\hat{F}_{m}(x)-F(x)|\leq\delta\right\}.$
Since $\hat{F}_{m}$ converges to $F$ uniformly in probability, we have $P(\mathscr{A}_{m})\xrightarrow{m\to\infty}1$.
On the event $\mathscr{A}_{m}$, for all $x\in{\bf R}$,
$|\hat{F}_{m}(x)-F(x)|\leq\delta.$
@@
$|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\leq\max\left\{|F^{-1}(u-\delta)-F^{-1}(u)|,\ |F^{-1}(u+\delta)-F^{-1}(u)|\right\}<\frac{\varepsilon}{2}.$
@@
$P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\geq\varepsilon\right)\leq P(A_{m}^{c})\xrightarrow{m\to\infty}0.$
@@
$\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
@@
$|u-v|<\delta\implies|F^{-1}(u)-F^{-1}(v)|<\frac{\varepsilon}{3}.$ (11)
@@
$\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\xrightarrow{P}0\quad\text{as }m\to\infty.$
@@
$P\left(\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|\geq\frac{\varepsilon}{3}\right)<\frac{\eta}{2}.$ (12)
@@
$\mathscr{B}_{m}=\left\{\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F^{-1}(u)|<\frac{\varepsilon}{3}\right\}.$
Then, $P(\mathscr{B}_{m})>1-\frac{\eta}{2}$ for all $m\geq M$.
@@
Here, the first and third terms are bounded by $\frac{\varepsilon}{3}$ due to (15), and the middle term is bounded by $\frac{\varepsilon}{3}$ due to the uniform continuity of $F^{-1}$ in (11).

### v3

@@ B.1 Crucial tools
Let $Q=\{Q_{i}\}_{i=1}^{m}$ be i.i.d. $d$-dimensional random vectors from a distribution $F$ on $\mathbb{R}^{d}$, where $Q_{i}=(Q_{i,1},\dots,Q_{i,d})$. Let $\hat{Q}=\{\hat{Q}_{i}\}_{i=1}^{m}$ be a perturbed version of $Q$, which can be potentially dependent. Let $K:\mathbb{R}^{d-1}\to[0,\infty)$ be a bounded (by a constant $K_{\max}$) and Lipschitz continuous kernel with $\int_{\mathbb{R}^{d-1}}K(u)\,\mathrm{d}u=1$. Let $h>0$ be a bandwidth parameter. Define the kernel-based conditional CDF estimator:
$\hat{F}_{m}^{d\,|\,1:(d-1)}(x\,|\,u):=\frac{\sum_{i=1}^{m}\mathbf{1}_{\{\hat{Q}_{i,d}\leq x\}}K\left(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\right)}{\sum_{i=1}^{m}K\left(\tfrac{u-\hat{Q}_{i,1:(d-1)}}{h}\right)},$
@@
  • The marginal density of the conditioning variables, $f_{1:(d-1)}(u)$, is continuous and bounded. Furthermore, there exists a compact set $\mathcal{U}\subset\mathbb{R}^{d-1}$ such that $\inf_{u\in\mathcal{U}}f_{1:(d-1)}(u)>0$.
  • The conditional CDF $F^{d\,|\,1:(d-1)}(x\,|\,u)$ is Lipschitz continuous with respect to $u=(x_{1},\dots,x_{d-1})$.
@@
$\sup_{u\in\mathcal{U},\,x\in\mathbb{R}}\big|\hat{F}_{m}^{d\,|\,1:(d-1)}(x\,|\,u)-F^{d\,|\,1:(d-1)}(x\,|\,u)\big|=o_{\mathbb{P}}(1).$
Hence $\hat{F}_{m}^{d\,|\,1:(d-1)}$ converges uniformly to $F^{d\,|\,1:(d-1)}$ in probability over $\mathcal{U}\times\mathbb{R}$.
@@
Let $d^{\prime}=d-1$. Define the oracle estimator $\tilde{F}_{m}^{d\,|\,1:d^{\prime}}(x\,|\,u)$ to be the same kernel estimator but constructed using the unperturbed data $\{Q_{i}\}$. By the triangle inequality, the error decomposes as:
$\sup_{u\in\mathcal{U},\,x\in\mathbb{R}}\big|\hat{F}_{m}^{d\,|\,1:d^{\prime}}(x\,|\,u)-F^{d\,|\,1:d^{\prime}}(x\,|\,u)\big|$ $\leq\underbrace{\sup_{u\in\mathcal{U},\,x\in\mathbb{R}}\big|\hat{F}_{m}^{d\,|\,1:d^{\prime}}(x\,|\,u)-\tilde{F}_{m}^{d\,|\,1:d^{\prime}}(x\,|\,u)\big|}_{(\textsf{I})}$
$\quad+\underbrace{\sup_{u\in\mathcal{U},\,x\in\mathbb{R}}\big|\tilde{F}_{m}^{d\,|\,1:d^{\prime}}(x\,|\,u)-F^{d\,|\,1:d^{\prime}}(x\,|\,u)\big|}_{(\textsf{II})}.$
Term (II): Standard Nonparametric Convergence.
Term (II) represents the uniform error of a standard kernel-based conditional CDF estimator. Under the assumptions that the marginal density $f_{1:d^{\prime}}$ is bounded away from zero on $\mathcal{U}$ and the conditional CDF is Lipschitz, standard results on ukernel estimators (e.g., from Wasserman [2006] for rates of kernel regression) provide the following rate:
$\sup_{u\in\mathcal{U},\,x\in\mathbb{R}}\big|\tilde{F}_{m}^{d\,|\,1:d^{\prime}}(x\,|\,u)-F^{d\,|\,1:d^{\prime}}(x\,|\,u)\big|=\mathcal{O}_{\mathbb{P}}\left(h+\sqrt{\frac{\log m}{mh^{d^{\prime}}}}\right).$
Here, $h$ corresponds to the smoothing bias and the second term corresponds to the uniform stochastic fluctuation (including the logarithmic factor for uniformity).
Term (I): Perturbation Analysis.
We analyze the effect of the data perturbation. For fixed $(x,u)$, write the estimators as ratios:
$\hat{F}_{m}^{d\,|\,1:d^{\prime}}(x\,|\,u)=\frac{A(x,u)}{B(u)},\quad\tilde{F}_{m}^{d\,|\,1:d^{\prime}}(x\,|\,u)=\frac{C(x,u)}{D(u)},$
@@
$A(x,u)=\sum_{i=1}^{m}\mathbf{1}_{\{\hat{Q}_{i,d}\leq x\}}K\biggl(\frac{u-\hat{Q}_{i,1:d^{\prime}}}{h}\biggr),\quad B(u)=\sum_{i=1}^{m}K\biggl(\frac{u-\hat{Q}_{i,1:d^{\prime}}}{h}\biggr),$
$C(x,u)=\sum_{i=1}^{m}\mathbf{1}_{\{Q_{i,d}\leq x\bigr\}}K\biggl(\frac{u-Q_{i,1:d^{\prime}}}{h}\biggr),\quad D(u)=\sum_{i=1}^{m}K\biggl(\frac{u-Q_{i,1:d^{\prime}}}{h}\biggr).$
The difference is bounded by:
$\left|\frac{A}{B}-\frac{C}{D}\right|=\left|\frac{(A-C)D-C(B-D)}{BD}\right|\leq\frac{|A-C|}{|B|}+\frac{|C|}{|B|}\frac{|B-D|}{|D|}.$
We proceed by establishing the stochastic orders of the components uniformly over $\mathcal{U}\times\mathbb{R}$.
  • Controlling $B$ and $D$: The sum $D(u)$ relates to the kernel density estimator $\hat{f}(u)$ via $D(u)=mh^{d^{\prime}}\hat{f}(u)$. Since $\hat{f}$ converges uniformly to $f_{1:d^{\prime}}$ and $\inf_{\mathcal{U}}f_{1:d^{\prime}}=c_{f}>0$, we have uniformly in probability:
    $D(u)\geq\frac{c_{f}}{2}mh^{d^{\prime}}\implies D(u)=\Omega_{\mathbb{P}}(mh^{d^{\prime}}).$
    Since $\delta_{m}\to 0$, we will show |BD||B-D| is of lower order, implying $B(u)=\Theta_{\mathbb{P}}(mh^{d^{\prime}})$ as well. Note also that $|C(x,u)|\leq D(u)$, so the ratio $|C/D|\leq 1$.
  • Controlling |BD||B-D|: Using the Lipschitz property of $K$ (with constant $L_{K}$) and the bound $\|\hat{Q}_{i}-Q_{i}\|\leq\delta_{m}$:
    $|B(u)-D(u)|\leq\sum_{i=1}^{m}\left|K\biggl(\frac{u-\hat{Q}_{i,1:d^{\prime}}}{h}\biggr)-K\biggl(\frac{u-Q_{i,1:d^{\prime}}}{h}\biggr)\right|.$
    The Lipschitz difference is bounded by $L_{K}\|\hat{Q}_{i,1:d^{\prime}}-Q_{i,1:d^{\prime}}\|/h\leq L_{K}\delta_{m}/h$. Importantly, this difference is non-zero only if the data point falls within the effective support of the kernel. Thhus this summation can be controlled in exactly the same manner as detailed in the proof of Lemma B.1. Briefly speaking, since $K$ has compact support, there exists a constant $R>0$ such that the support is contained within a hypercube $R,R]^{d^{\prime}}$.
    Consequently, the term for index $i$ is non-zero only if the unperturbed point $Q_{i,1:d^{\prime}}$ falls within a neighborhood of $u$ with radius roughly $Rh+\delta_{m}$. Since $\delta_{m}=o(h)$, this effective support is contained within a hypercube $\mathcal{C}_{u,h}$ centered at $u$ with side length proportional to $h$. The volume of this hypercube is $\mathcal{O}(h^{d^{\prime}})$.
    Let $N_{h}(u)$ be the number of data points $Q_{i,1:d^{\prime}}$ falling into this local neighborhood. Since the marginal density $f_{1:d^{\prime}}$ is bounded by $C_{\mathrm{up}}$, the probability that a single observation falls into this region is bounded by:
    $\mathbb{P}(Q_{i,1:d^{\prime}}\in\mathcal{C}_{u,h})\leq C_{\mathrm{up}}\cdot\mathsf{Vol}(\mathcal{C}_{u,h})=\mathcal{O}(h^{d^{\prime}}).$
    Thus, the number of active points $N_{h}(u)$ behaves as a sum of Bernoulli trials with success probability $\mathcal{O}(h^{d^{\prime}})$. The expected count is $\mathcal{O}(mh^{d^{\prime}})$. Since the density is bounded, the count is uniformly bounded in probability by the same order, i.e., $\sup_{u}N_{h}(u)=\mathcal{O}_{\mathbb{P}}(mh^{d^{\prime}})$.
    Summing the Lipschitz error bound over these active points yields:
    $\sup_{u\in\mathcal{U}}|B(u)-D(u)|\leq N_{h}(u)\frac{L_{K}\delta_{m}}{h}=\mathcal{O}_{\mathbb{P}}\left(mh^{d^{\prime}}\frac{\delta_{m}}{h}\right)=\mathcal{O}_{\mathbb{P}}(mh^{d^{\prime}-1}\delta_{m}).$
    When we have $\delta_{m}=o(h)$, this term is in a lower order compared with $D(u)$ and thus $B(u)$ is the same order of $D(u)$.
  • Controlling |AC||A-C|: We decompose the numerator difference:
    $|A-C|$ $=\left|\sum_{i=1}^{m}\mathbf{1}_{\{\hat{Q}_{i,d}\leq x\}}K\biggl(\frac{u-\hat{Q}_{i,1:d^{\prime}}}{h}\biggr)-\sum_{i=1}^{m}\mathbf{1}_{\{Q_{i,d}\leq x\}}K\biggl(\frac{u-Q_{i,1:d^{\prime}}}{h}\biggr)\right|$
    $\leq\sum_{i=1}^{m}\mathbf{1}_{\{\hat{Q}_{i,d}\leq x\}}\left|K\biggl(\frac{u-\hat{Q}_{i,1:d^{\prime}}}{h}\biggr)-K\biggl(\frac{u-Q_{i,1:d^{\prime}}}{h}\biggr)\right|$
    $\quad+\sum_{i=1}^{m}\big|\mathbf{1}_{\{\hat{Q}_{i,d}\leq x\}}-\mathbf{1}_{\{Q_{i,d}\leq x\}}\big|K\biggl(\frac{u-Q_{i,1:d^{\prime}}}{h}\biggr)$
    $=\Delta_{\mathrm{kernel}}+\Delta_{\mathrm{ind}}.$
    $\Delta_{\mathrm{kernel}}$ can be directly bounded by |BD||B-D|, since the indicator is $\leq 1$. Thus,
    $\Delta_{\mathrm{kernel}}=\mathcal{O}_{\mathbb{P}}(mh^{d^{\prime}-1}\delta_{m}).$
    Combining and merging the bounds, we have
    $\sup_{u\in\mathcal{U},\,x\in\mathbb{R}}|A(x,u)-C(x,u)|=\mathcal{O}_{\mathbb{P}}(mh^{d^{\prime}-1}\delta_{m}).$
Substituting these rates back into Term (I), we get
@@
$\frac{|A-C|}{|B|}=\frac{\mathcal{O}_{\mathbb{P}}(mh^{d^{\prime}-1}\delta_{m})}{\Theta_{\mathbb{P}}(mh^{d^{\prime}})}=\mathcal{O}_{\mathbb{P}}\left(\frac{\delta_{m}}{h}\right),$
$\frac{|B-D|}{|D|}=\frac{\mathcal{O}_{\mathbb{P}}(mh^{d^{\prime}-1}\delta_{m})}{\Theta_{\mathbb{P}}(mh^{d^{\prime}})}=\mathcal{O}_{\mathbb{P}}\left(\frac{\delta_{m}}{h}\right).$
Therefore, we have Term (I) to be $\mathcal{O}_{\mathbb{P}}(\delta_{m}/h)$.
Combining Term (I) and Term (II) yields
@@
$\sup_{u\in\mathcal{U},\,x\in\mathbb{R}}\big|\hat{F}_{m}^{d\,|\,1:d^{\prime}}(x\,|\,u)-F^{d\,|\,1:d^{\prime}}(x\,|\,u)\big|=\mathcal{O}_{\mathbb{P}}\left(h+\sqrt{\frac{\log m}{mh^{d-1}}}+\frac{\delta_{m}}{h}\right).$
@@
Let $F$ be a continuous and strictly increasing cumulative distribution function (CDF) (or conditional CDF) on $\mathbb{R}$, and let $F^{-1}$ denote its inverse function. Assume the conditions of Corollary B.1 hold. Let $\hat{F}_{m}$ be the smoothed CDF estimator of Corollary B.1. Then, for any closed interval $a,b]\subset(0,1)$, the inverse functions $\hat{F}_{m}^{-1}$ converge uniformly in probability to $F^{-1}$ on $[a,b]$; that is
$\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m\to\infty.$
@@
$\mathbb{P}\left(\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta\end{subarray}}\big|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)\big|\geq\varepsilon\right)<\eta.$
@@
$\mathbb{P}\left(\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|>\varepsilon\right)\xrightarrow{m\to\infty}0.$
@@
$|u-v|<\delta\implies\big|F^{-1}(u)-F^{-1}(v)\big|<\frac{\varepsilon}{2}.$
@@
$\mathscr{A}_{m}=\left\{\sup_{x\in\mathbb{R}}\big|\hat{F}_{m}(x)-F(x)\big|\leq\delta\right\}.$
Since $\hat{F}_{m}$ converges to $F$ uniformly in probability, we have $\mathbb{P}(\mathscr{A}_{m})\xrightarrow{m\to\infty}1$.
On the event $\mathscr{A}_{m}$, for all $x\in\mathbb{R}$,
$\big|\hat{F}_{m}(x)-F(x)\big|\leq\delta.$
@@
$\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|\leq\max\left\{\big|F^{-1}(u-\delta)-F^{-1}(u)\big|,\,\big|F^{-1}(u+\delta)-F^{-1}(u)\big|\right\}<\frac{\varepsilon}{2}.$
@@
$\mathbb{P}\left(\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|\geq\varepsilon\right)\leq\mathbb{P}(\mathscr{A}_{m}^{c})\xrightarrow{m\to\infty}0.$
@@
$\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m\to\infty.$
@@
$|u-v|<\delta\implies\big|F^{-1}(u)-F^{-1}(v)\big|<\frac{\varepsilon}{3}.$ (11)
@@
$\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m\to\infty.$
@@
$\mathbb{P}\left(\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|\geq\frac{\varepsilon}{3}\right)<\frac{\eta}{2}.$ (12)
@@
$\mathscr{B}_{m}=\left\{\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F^{-1}(u)\big|<\frac{\varepsilon}{3}\right\}.$
Then, $\mathbb{P}(\mathscr{B}_{m})>1-\eta/2$ for all $m\geq M$.
@@
Here, the first and third terms are bounded by $\varepsilon/3$ due to (12), and the middle term is bounded by $\varepsilon/3$ due to the uniform continuity of $F^{-1}$ in (11).